A classifier of fake banknotes is being tested to see how reliable it can detect real banknotes from fake ones. Three confusion matrices representing 3 Days of testing was made


Day 1

n = 100

Predicted No Predicted Yes

Actual No - 15 35

Actual Yes - 8 42



Day 2

n = 100

Predicted No Predicted Yes

Actual No - 22 28

Actual Yes - 2 48



Day 3

n = 100


Predicted No Predicted Yes

Actual No - 43 7

Actual Yes - 16 34


1) What Day has the best accuracy

A) Day 1

B) Day 2

C) Day 3

D) Insufficient Information. Cannot be determined


2)How many tests are there in total?

3)What is the F-Score of Day 3?




1
Expert's answer
2021-05-20T14:40:56-0400

1) Day 1 has the best accuracy due to the large value of Actual yes ratio as compared to the other days.


2) There are total 3 test conducted on 3 days. as the activity are carried out in 3 days. So we require three test .


3) F-score of Day-3 is

"F=\dfrac{TP}{TP+\frac{1}{2}(FP+FN)}"


Where, TP=True positives

    FP=False positives

    FN=False negatives


"f=\dfrac{1634}{1634+\frac{1}{2}(437+1197)}\\[9pt]=\dfrac{1634}{1634+817}\\[9pt]=\dfrac{1634}{2451}=0.66"